A shopkeeper increased the price of a cycle by 15%. The increased price was Rs. 1, 610. What was the original price
Answer by Maths68(1474)

Let
Original Price of Cycle = x
15% of x = x*0.15=0.15x
Given
Increased price = 1610
Original Price + 15% of x = Increased price
x+0.15x=1610
1.15x=1610
1.15x/1.15=1610/1.15
x=1400
Original Price of Cycle = Rs. 1400